# Global analysis of charm mixing parameters and determination of the CKM angle $\mathbf{\gamma}$

### v1

@@ VI Conclusions
We presented a combination of charm and beauty observables to determine the CKM angle $\gamma$ and the neutral $D$ mixing and CP-violating parameters simultaneously in the framework of the SM and employing approximate universality for $D^{0}-\overline{D^{0}}$ mixing.
@@
$x_{12}\simeq x=(0.402\pm 0.034)\%,\qquad y_{12}\simeq y=(0.629\pm 0.021)\%,\qquad\qquad\phi_{2}^{M}=(2.4\pm 1.6)\degree,\qquad\phi_{2}^{\Gamma}=(2.6\pm 1.2)\degree,$
$\phi_{2}=(-2.54\pm 0.94)\degree,\qquad\qquad\left|q/p\right|-1=(-0.1\pm 1.6)\%,$
@@
$\gamma=(68.4\pm 2.6)\degree.$
The uncertainties on the two CP-violating phases $\phi_{2}^{M,\Gamma}$ are still one order of magnitude larger than what is needed to test the predictions $\phi_{2}^{M}\sim\phi_{2}^{\Gamma}\sim 0.1\degree$, obtained from U-spin arguments in Eq. (10), but a factor of four larger than the upper bound of $\left|\phi_{2}^{\Gamma}\right|<0.3\degree$ in Eq. (11). An interesting byproduct of our fit is the value of the direct CP asymmetries in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ decays within the framework of approximate universality:
$a_{D}^{KK}=(8.5\pm 6.3)\cdot 10^{-4}\,,\qquad a_{D}^{\pi\pi}=(30.1\pm 7.0)\cdot 10^{-4}\,,\qquad\rho=67\%\,,$ (45)
providing evidence for direct CP violation in $D^{0}\to\pi^{+}\pi^{-}$ at more than $4\sigma$, together with a sizable deviation from the U-spin expectation of $a_{D}^{KK}\sim-a_{D}^{\pi\pi}$.
@@
$\gamma_{B^{\pm}}=(67.4\pm 3.1)\degree,\qquad\qquad\gamma_{B^{0}}=(69.5\pm 5.7)\degree,\qquad\qquad\gamma_{B^{0}_{s}}=(76.8\pm 9.6)\degree.$
@@ Acknowledgements.
We warmly thank F. Blanc, A. Gilman, M. Morello and S. Vecchi for correspondence on LHCb results. This work was supported in part by the European Union - Next Generation EU under italian MUR grant PRIN-2022-RXEZCJ and by the HSE basic research fund.
@@ Appendix A GLW Observables
$A_{B}^{\text{CP}}(f,h)=\frac{\Gamma(B\to[f]_{D}h)-\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})}=\frac{2r_{B}^{Dh}\kappa_{B}^{Dh}\sin\gamma\sin\delta_{B}^{Dh}}{(1+(r_{B}^{Dh})^{2})\frac{1-(2F_{+}^{f}-1)\alpha y}{2F_{+}^{f}-1-\alpha y}+2r_{B}^{Dh}\kappa_{B}^{Dh}\cos\gamma\cos\delta_{B}^{Dh}}.$ (46)
@@
$R(f,f^{*},h)=\frac{\mathcal{B}(D^{0}\to f^{*})}{\mathcal{B}(D^{0}\to f)}\frac{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[f^{*}]_{D}h)+\Gamma(\overline{B}\to[\overline{f^{*}}]_{D}\overline{h})}.$ (47)
@@ Appendix B ADS Observables
$R^{-}(f,h)=\frac{\Gamma(B\to[\overline{f}]_{D}h)}{\Gamma(B\to[f]_{D}h)},\qquad\qquad R^{+}(f,h)=\frac{\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})}.$ (51)
@@
$\Gamma(B\to[\overline{f}]_{D}h)\propto(r_{D}^{f})^{2}+(r_{B}^{Dh})^{2}+2r_{B}^{Dh}r_{D}^{f}\kappa_{B}^{Dh}\kappa_{D}^{f}\cos(\delta_{B}^{Dh}+\delta_{D}^{f}-\gamma)+\Gamma_{\mathrm{mix}}(-x,y,\gamma,f,h),$ (52)
@@
$A^{\text{sup}}(f,h)=\frac{\Gamma(B\to[\overline{f}]_{D}h)-\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[\overline{f}]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})},\qquad\qquad A^{\text{fav}}(f,h)=\frac{\Gamma(B\to[f]_{D}h)-\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})}{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})}.$ (53)
@@
$R^{\text{ADS}}(f,h)=\frac{\Gamma(B\to[\overline{f}]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})},$ (57)
or using two different channels for beauty decays, $B\to[f]_{D}h$ and $B\to[f]_{D}h^{\prime}$, through the following observables
$R^{\text{sup}}(f,h,h^{\prime})=\frac{\Gamma(B\to[\overline{f}]_{D}h)+\Gamma(\overline{B}\to[f]_{D}\overline{h})}{\Gamma(B\to[\overline{f}]_{D}h^{\prime})+\Gamma(\overline{B}\to[f]_{D}\overline{h^{\prime}})},\qquad\qquad R^{\text{fav}}(f,h,h^{\prime})=\frac{\Gamma(B\to[f]_{D}h)+\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h})}{\Gamma(B\to[f]_{D}h^{\prime})+\Gamma(\overline{B}\to[\overline{f}]_{D}\overline{h^{\prime}})}.$ (58)
@@
$R^{\text{sup}}(f,h,h^{\prime})=\frac{\mathcal{B}(B\to D^{0}h)}{\mathcal{B}(B\to D^{0}h^{\prime})}\frac{(r_{B}^{Dh})^{2}+(r_{D}^{f})^{2}+2r_{D}^{f}r_{B}^{Dh}\kappa_{D}^{f}\kappa_{B}^{Dh}\cos\gamma\cos(\delta_{B}^{Dh}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(-x,y,\gamma,f,h)}{(r_{B}^{Dh^{\prime}})^{2}+(r_{D}^{f})^{2}+2r_{D}^{f}r_{B}^{Dh^{\prime}}\kappa_{D}^{f}\kappa_{B}^{Dh^{\prime}}\cos\gamma\cos(\delta_{B}^{Dh^{\prime}}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(-x,y,\gamma,f,h^{\prime})},$ (60)
@@
$R^{\text{fav}}(f,h,h^{\prime})=\frac{\mathcal{B}(B\to D^{0}h)}{\mathcal{B}(B\to D^{0}h^{\prime})}\frac{1+(r_{D}^{f}r_{B}^{Dh})^{2}+2r_{B}^{Dh}\kappa_{B}^{Dh}r_{D}^{f}\kappa_{D}^{f}\cos\gamma\cos(\delta_{B}^{Dh}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h)}{1+(r_{D}^{f}r_{B}^{Dh^{\prime}})^{2}+2r_{B}^{Dh^{\prime}}\kappa_{B}^{Dh^{\prime}}r_{D}^{f}\kappa_{D}^{f}\cos\gamma\cos(\delta_{B}^{Dh^{\prime}}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h^{\prime})}.$ (61)
Appendix C Parameters of the combination
Parameter Range Parameter Range Parameter Range
$\phi_{2}^{M}[\degree]$ $[-11,17]$ ϕ2Γ[°]superscriptsubscriptitalic-$\phi_{2}^{\Gamma}[\degree]$ $[-11,11]$ $F_{+}^{4\pi}$ $[0.5,1]$
$x_{12}\simeq x[\%]$ $[0.05,0.8]$ $y_{12}\simeq y[\%]$ $[0.3,0.9]$ $F_{+}^{\pi\pi\pi^{0}}$ $[0.6,1]$
$a_{D}^{KK}$ $[-0.1,0.1]$ $a_{D}^{\pi\pi}$ $[-0.1,0.1]$ $\Delta Y^{KK}-\Delta Y^{\pi\pi}$ $[-0.1,0.1]$
$(\left\langle\tau\right\rangle^{KK}_{E^{\prime},\pi}+\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime},\pi})/2$ $[1,3]$ $(\left\langle\tau\right\rangle^{KK}_{E^{\prime},\mu}+\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime},\mu})/2$ $[1.1,1.3]$ $\left\langle\tau\right\rangle^{KK}_{E^{\prime},\pi}-\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime},\pi}$ $[0.12,0.15]$
$\left\langle\tau\right\rangle^{KK}_{E^{\prime},\mu}-\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime},\mu}$ $[-0.01,0.01]$ $\left\langle\tau\right\rangle^{KK}_{E,C_{D^{+}}}$ $[1.46,1.8]$ $\left\langle\tau\right\rangle^{KK}_{E,C_{D_{s}}}$ $[1.6,1.8]$
$\kappa_{D}^{K^{0}_{S}K\pi}$ $[0.3,1]$ $\delta_{D}^{K^{0}_{S}K\pi}[\degree]$ $[-115,115]$ $F_{+}^{KK\pi^{0}}$ $[0.2,1]$
$r_{D}^{K\pi}$ $[0.055,0.065]$ $\delta_{D}^{K\pi}[\degree]$ $[115,258]$ $F_{+}^{KK\pi\pi}$ $[0,1]$
$r_{D}^{K3\pi}$ $[0.048,0.065]$ $\delta_{D}^{K3\pi}[\degree]$ $[29,286]$ $\kappa_{D}^{K3\pi}$ $[0,1]$
$r_{D}^{K\pi\pi^{0}}$ $[0.03,0.06]$ $\delta_{D}^{K\pi\pi^{0}}[\degree]$ $[86,275]$ $\kappa_{D}^{K\pi\pi^{0}}$ $[0.4,1]$
$r_{D}^{K^{0}_{S}K\pi}$ $[0.55,0.7]$
Table 7: Ranges of the uniform priors used for the parameters describing the charm part of the combination. The notation of Sec. III is adopted. Notice that the subscripts $\pi$ and $\mu$ in $\left\langle\tau\right\rangle^{f}_{E^{\prime},\pi}$ and $\left\langle\tau\right\rangle^{f}_{E^{\prime},\mu}$ refer to two different ways of tagging the neutral $D$ mesons. More information are available in ref. Aaij et al. (2019a). The subscripts $C_{D^{+}}$ and $C_{D_{s}}$ in $\left\langle\tau\right\rangle^{f}_{E,D^{+}}$ and $\left\langle\tau\right\rangle^{f}_{E,D_{s}}$ refers to the two strategies adopted in ref. Aaij et al. (2023a) for background subtraction.
Parameter Range Parameter Range Parameter Range Parameter Range
$r_{B}^{DK}$ $[0.07,0.13]$ $r_{B}^{D\pi}$ $[0,0.012]$ $\delta_{B}^{DK}[\degree]$ $[86,172]$ $\delta_{B}^{D\pi}[\degree]$ $[172,360]$
$\mathbf{B^{\pm}}$ $r_{B}^{D^{*}K}$ $[0,0.3]$ $r_{B}^{D^{*}\pi}$ $[0,0.06]$ $\delta_{B}^{D^{*}K}[\degree]$ $[-229,57]$ $\delta_{B}^{D^{*}\pi}[\degree]$ $[-57,303]$
Parameters $r_{B}^{DK^{*\pm}}$ $[0,0.3]$ $\delta_{B}^{DK^{*\pm}}[\degree]$ $[-86,258]$ $\kappa_{B}^{DK^{*\pm}}$ $[0.5,1]$ $r_{B}^{DK\pi\pi}$ $[0,0.2]$
$r_{B}^{D\pi\pi\pi}$ $[0,0.2]$ $\delta_{B}^{DK\pi\pi}[\degree]$ $[-266,94]$ $\delta_{B}^{D\pi\pi\pi}[\degree]$ $[-266,94]$ $\kappa_{B}^{DK\pi\pi}$ $[0,1]$
$\kappa_{B}^{D\pi\pi\pi}$ $[0,1]$ $\frac{\mathcal{B}(B\to DK)}{\mathcal{B}(B\to D\pi)}$ $[0.05,0.1]$
CKM Angles $\gamma[\degree]$ $[40,86]$ $\beta[\degree]$ $[11,29]$ $-2\beta_{s}[\degree]$ $[-17,11]$
$\mathbf{B^{0}}$ $r_{B^{0}}^{DK^{*0}}$ $[0,0.5]$ $\delta_{B^{0}}^{DK^{*0}}[\degree]$ $[115,315]$ $\kappa_{B^{0}}^{DK^{*0}}$ $[0.75,1]$ $r_{B^{0}}^{D\pi}$ $[0,0.4]$
Parameters $\delta_{B^{0}}^{D\pi}[\degree]$ $[-180,180]$
$\mathbf{B^{0}_{s}}$ $r_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $[0,0.5]$ $\delta_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $[-105,255]$ $\kappa_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $[0,1]$ $\delta_{B^{0}_{s}}^{D_{s}K\pi\pi}[\degree]$ $[-115,86]$
Parameters $\kappa_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $[0,1]$ $r_{B^{0}_{s}}^{D_{s}K}$ $[0,1.2]$ $\delta_{B^{0}_{s}}^{D_{s}K}[\degree]$ $[-143,143]$ $r_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $[0,1.2]$
Table 8: Ranges of the uniform priors employed for the parameters describing the beauty observables, divided by the species of the $B$ mesons involved. The notation of Sec. IV is adopted.
We report the intervals defining the uniform priors adopted for the charm parameters in Tab. 7, while the beauty parameters are contained in Tab. 8, divided by types of $B$ mesons involved in the observables. The intervals are chosen according to the definitions of the parameters and to speed up the convergence of the Markov chains, after having checked that no solutions are lost.
Appendix D Fit results
@@
Figure 6: Two-dimensional contours of the charm and beauty decay parameters describing the most precise observables of the combination and the direct CP asymmetries in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ decays. Darker (lighter) regions correspond to $68.3\%$ ($95.4\%$) probability.
$\gamma$ $x_{12}$ $y_{12}$ $\phi_{2}^{M}$ $\phi_{2}^{\Gamma}$ $\phi_{12}$ $r_{B}^{DK}$ $\delta_{B}^{DK}$ $r_{B}^{D\pi}$ $\delta_{B}^{D\pi}$ $r_{D}^{K\pi}$ $\delta_{D}^{K\pi}$ $\phi_{2}$ $|q/p|-1$ $a_{D}^{KK}$ $a_{D}^{\pi\pi}$
$\gamma$ $100$- $1$- -- -- $43$ $50$ $14$ $24$- $4$- -- -- --
$x_{12}$ $100$ $4$ $-12$ $1$ $-10$ $1$ $2$ $-2$ $-1$ $8$ $2$ $5$ $-10$ $1$-
$y_{12}$ $100$ $-1$ $-6$ $2$ $-10$ $-20$ $4$ $-2$ $-14$ $42$ $6$ $2$- --
$\phi_{2}^{M}$ $100$ $-4$ $81$- $1$- -- $-3$ $-3$ $-44$ $81$ $31$ $26$
$\phi_{2}^{\Gamma}$ $100$ $-62$ $1$ $3$- $1$- $-5$ $-87$ $-62$ $-1$ $-1$
$\phi_{12}$ $100$- $-1$- -- $-3$ $1$ $17$ $100$ $24$ $21$
$r_{B}^{DK}$ $100$ $33$ $4$ $11$ $-9$ $-21$ $-1$- -- --
$\delta_{B}^{DK}$ $100$ $8$ $23$ $-14$ $-51$ $-3$ $-1$- --
$r_{B}^{D\pi}$ $100$ $68$ $13$- -- -- -- --
$\delta_{B}^{D\pi}$ $100$ $11$ $-18$- -- -- --
$r_{D}^{K\pi}$ $100$ $28$ $1$ $-3$- --
$\delta_{D}^{K\pi}$ $100$ $6$ $1$- --
$\phi_{2}$ $100$ $17$ $-14$ $-12$
$|q/p|-1$ $100$ $24$ $21$
$a_{D}^{KK}$ $100$ $67$
@@
Table 9: Correlation matrix for the most significant quantities in percent. Values smaller than $1\%$ are not reported. Large correlations between $\phi_{2}^{M,\Gamma}$, $\phi_{12}$, $|q/p|-1$ and $\phi_{2}$ are due to the fact that they are not independent parameters, as shown in Eqs. (7) and (13).
We report the results of the fit obtained by combining all the observables.
The correlation matrix of the most physically relevant parameters is reported in Tab. 9.
Tabs. 10 and 11 present the probability intervals found for the charm and beauty parameters, respectively. In Fig. 6, we report two-dimensional contours of the ratio of decay amplitudes and strong phases describing the most precise modes used in the combination: $B^{\pm}\to D\pi^{\pm}(K^{\pm})$ and $D\to K^{\pm}\pi^{\mp}$, together with the direct CP asymmetries between the charm decay amplitudes of the modes $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$.

### v3

@@ VI Conclusions
We presented a Bayesian combination of charm and beauty observables to determine the CKM angle $\gamma$ and the neutral $D$ mixing and CP-violating parameters simultaneously in the framework of the SM and employing approximate universality for $D^{0}-\overline{D^{0}}$ mixing.
@@
$x_{12}\simeq x=(0.401\pm 0.043)\%,\qquad y_{12}\simeq y=(0.610\pm 0.017)\%,\qquad\qquad\phi_{2}^{M}=(0.13\pm 0.70)\degree,\qquad\phi_{2}^{\Gamma}=(2.1\pm 1.6)\degree,$
$\phi_{2}=(-1.5\pm 1.1)\degree,\qquad\qquad\left|q/p\right|-1=(-1.6\pm 1.5)\%,$
@@
$\gamma=(65.7\pm 2.5)\degree,$
in perfect agreement with the indirect determination from the Unitary triangle analysis $\gamma_{\mathrm{UT}}=(65.6\pm 1.4)\degree$ ICHEP24-Bona . The uncertainties on the two CP-violating phases $\phi_{2}^{M,\Gamma}$ are still one order of magnitude larger than what is needed to test the predictions $\phi_{2}^{M}\sim\phi_{2}^{\Gamma}\sim 0.1\degree$ obtained from U-spin arguments in Eq. (10), and a factor of seven larger than the upper bound of $\left|\phi_{2}^{\Gamma}\right|<0.3\degree$ in Eq. (11). The direct CP asymmetries in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ decays within the framework of approximate universality are given by
$a_{D}^{KK}=(4.0\pm 5.3)\cdot 10^{-4}\,,\qquad a_{D}^{\pi\pi}=(23.4\pm 6.0)\cdot 10^{-4}\,.$ (46)
@@
$\gamma_{B^{\pm}}=(65.5\pm 2.7)\degree,\qquad\qquad\gamma_{B^{0}}=(57\pm 13)\degree,\qquad\qquad\gamma_{B^{0}_{s}}=(77.4\pm 9.6)\degree.$
@@ Acknowledgements.
We warmly thank F. Blanc, A. Gilman, M. Morello and S. Vecchi for correspondence on LHCb results. We are indebted to T. Pajero, M.W. Kenzie, A. Gilman, J. Butter and M.P. Whitehead for pointing out to us several problems in the first version of this manuscript. This work was supported in part by the European Union - Next Generation EU under italian MUR grant PRIN-2022-RXEZCJ and by the HSE basic research fund.
@@ Appendix A GLW Observables
$A_{B}^{\text{CP}}(f,h)=\frac{\Gamma(\overline{B}\to[f]_{D}h)-\Gamma(B\to[f]_{D}\overline{h})}{\Gamma(\overline{B}\to[f]_{D}h)+\Gamma(B\to[f]_{D}\overline{h})}=\frac{2r_{B}^{Dh}\kappa_{B}^{Dh}\sin\gamma\sin\delta_{B}^{Dh}}{(1+(r_{B}^{Dh})^{2})\frac{1-(2F_{+}^{f}-1)\alpha y}{2F_{+}^{f}-1-\alpha y}+2r_{B}^{Dh}\kappa_{B}^{Dh}\cos\gamma\cos\delta_{B}^{Dh}}.$ (47)
@@
$R(f,f^{*},h)=\frac{\mathcal{B}(D^{0}\to f^{*})}{\mathcal{B}(D^{0}\to f)}\frac{\Gamma(\overline{B}\to[f]_{D}h)+\Gamma(B\to[f]_{D}\overline{h})}{\Gamma(\overline{B}\to[f^{*}]_{D}h)+\Gamma(B\to[\overline{f^{*}}]_{D}\overline{h})}.$ (48)
@@ Appendix B ADS Observables
$R^{-}(f,h)=\frac{\Gamma(\overline{B}\to[\overline{f}]_{D}h)}{\Gamma(\overline{B}\to[f]_{D}h)},\qquad\qquad R^{+}(f,h)=\frac{\Gamma(B\to[f]_{D}\overline{h})}{\Gamma(B\to[\overline{f}]_{D}\overline{h})}.$ (52)
@@
$\Gamma(\overline{B}\to[\overline{f}]_{D}h)\propto(r_{D}^{f})^{2}+(r_{B}^{Dh})^{2}+2r_{B}^{Dh}r_{D}^{f}\kappa_{B}^{Dh}\kappa_{D}^{f}\cos(\delta_{B}^{Dh}+\delta_{D}^{f}-\gamma)+\Gamma_{\mathrm{mix}}(-x,y,\gamma,f,h),$ (53)
@@
$A^{\text{sup}}(f,h)=\frac{\Gamma(\overline{B}\to[\overline{f}]_{D}h)-\Gamma(B\to[f]_{D}\overline{h})}{\Gamma(\overline{B}\to[\overline{f}]_{D}h)+\Gamma(B\to[f]_{D}\overline{h})},\qquad\qquad A^{\text{fav}}(f,h)=\frac{\Gamma(\overline{B}\to[f]_{D}h)-\Gamma(B\to[\overline{f}]_{D}\overline{h})}{\Gamma(\overline{B}\to[f]_{D}h)+\Gamma(B\to[\overline{f}]_{D}\overline{h})}.$ (54)
@@
$R^{\text{ADS}}(f,h)=\frac{\Gamma(\overline{B}\to[\overline{f}]_{D}h)+\Gamma(B\to[f]_{D}\overline{h})}{\Gamma(\overline{B}\to[f]_{D}h)+\Gamma(B\to[\overline{f}]_{D}\overline{h})},$ (58)
or using two different channels for beauty decays, $\overline{B}\to[f]_{D}h$ and $\overline{B}\to[f]_{D}h^{\prime}$, through the following observables
$R^{\text{sup}}(f,h,h^{\prime})=\frac{\Gamma(\overline{B}\to[\overline{f}]_{D}h)+\Gamma(B\to[f]_{D}\overline{h})}{\Gamma(\overline{B}\to[\overline{f}]_{D}h^{\prime})+\Gamma(B\to[f]_{D}\overline{h^{\prime}})},\qquad\qquad R^{\text{fav}}(f,h,h^{\prime})=\frac{\Gamma(\overline{B}\to[f]_{D}h)+\Gamma(B\to[\overline{f}]_{D}\overline{h})}{\Gamma(\overline{B}\to[f]_{D}h^{\prime})+\Gamma(B\to[\overline{f}]_{D}\overline{h^{\prime}})}.$ (59)
@@
$R^{\text{sup}}(f,h,h^{\prime})=\frac{\mathcal{B}(\overline{B}\to D^{0}h)}{\mathcal{B}(\overline{B}\to D^{0}h^{\prime})}\frac{(r_{B}^{Dh})^{2}+(r_{D}^{f})^{2}+2r_{D}^{f}r_{B}^{Dh}\kappa_{D}^{f}\kappa_{B}^{Dh}\cos\gamma\cos(\delta_{B}^{Dh}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(-x,y,\gamma,f,h)}{(r_{B}^{Dh^{\prime}})^{2}+(r_{D}^{f})^{2}+2r_{D}^{f}r_{B}^{Dh^{\prime}}\kappa_{D}^{f}\kappa_{B}^{Dh^{\prime}}\cos\gamma\cos(\delta_{B}^{Dh^{\prime}}+\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(-x,y,\gamma,f,h^{\prime})},$ (61)
@@
$R^{\text{fav}}(f,h,h^{\prime})=\frac{\mathcal{B}(\overline{B}\to D^{0}h)}{\mathcal{B}(\overline{B}\to D^{0}h^{\prime})}\frac{1+(r_{D}^{f}r_{B}^{Dh})^{2}+2r_{B}^{Dh}\kappa_{B}^{Dh}r_{D}^{f}\kappa_{D}^{f}\cos\gamma\cos(\delta_{B}^{Dh}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h)}{1+(r_{D}^{f}r_{B}^{Dh^{\prime}})^{2}+2r_{B}^{Dh^{\prime}}\kappa_{B}^{Dh^{\prime}}r_{D}^{f}\kappa_{D}^{f}\cos\gamma\cos(\delta_{B}^{Dh^{\prime}}-\delta_{D}^{f})+\Gamma_{\mathrm{mix}}^{+}(x,y,\gamma,f,h^{\prime})}.$ (62)
Appendix C Fit results
@@
Figure 8: Two-dimensional contours of the beauty decay parameters describing the most precise observables of the combination obtained by combining all the modes.
We report the results of the fit obtained by combining all the observables. The correlation matrix of the most physically relevant parameters is reported in Tab. 7. Tabs. 8,  9, and 10 present the probability intervals found for the charm and beauty parameters. In Fig. 8, we report two-dimensional contours of the ratio of decay amplitudes and strong phases describing the most precise modes used in the combination: $B^{\pm}\to D\pi^{\pm}(K^{\pm})$.
$\gamma$ $r_{B}^{DK}$ $\delta_{B}^{DK}$ $r_{B}^{D\pi}$ $\delta_{B}^{D\pi}$ $r_{D}^{K\pi}$ $\delta_{D}^{K\pi}$ $x_{12}$ $y_{12}$ $\phi_{2}^{M}$ $\phi_{2}^{\Gamma}$ $\phi_{2}$ $|q/p|-1$ $a_{D}^{KK}$ $a_{D}^{\pi\pi}$
$\gamma$ $100$ $35$ $58$ $3$ $18$ $1$ $5$- $5$ $-1$ $-1$ $1$- -- --
$r_{B}^{DK}$ $100$ $17$ $-6$ $2$ $-4$ $-9$ $1$ $-8$ $1$ $1$ $-1$ $-1$- --
$\delta_{B}^{DK}$ $100$ $-11$ $12$ $-14$ $-38$ $2$ $-33$ $4$ $5$ $-5$ $-3$- --
$r_{B}^{D\pi}$ $100$ $60$ $7$ $10$ $-1$ $8$ $-1$ $-1$ $1$ $1$- --
$\delta_{B}^{D\pi}$ $100$ $1$ $-9$ $-1$ $-10$ $1$ $2$ $-2$ $-1$- --
$r_{D}^{K\pi}$ $100$ $38$ $36$ $3$ $-5$ $-3$ $7$ $-1$ $-1$ $-1$
$\delta_{D}^{K\pi}$ $100$ $6$ $88$ $-11$ $-13$ $14$ $8$ $-1$ $-1$
$x_{12}$ $100$ $8$ $1$ $-1$ $8$ $-4$ $1$ $1$
$y_{12}$ $100$ $-10$ $-10$ $11$ $6$ $-1$-
$\phi_{2}^{M}$ $100$ $-18$ $1$ $53$ $43$ $42$
$\phi_{2}^{\Gamma}$ $100$ $-98$ $-93$ $3$ $1$
$\phi_{2}$ $100$ $84$ $-11$ $-9$
$|q/p|-1$ $100$ $14$ $15$
$a_{D}^{KK}$ $100$ $57$
@@
Table 7: Correlation matrix for the most significant quantities in percent. Values smaller than $1\%$ are not reported. Large correlations between $\phi_{2}^{M,\Gamma}$, $|q/p|-1$ and $\phi_{2}$ are due to the fact that they are not independent parameters, as shown in Eqs. (7) and (13).